# Automated Planning Domain Inference for Task and Motion Planning

### v2

@@ IV-C Precondition and Postcondition Generation
The precondition is determined by identifying the intersections of the pre-images of the same action across multiple instances. We first extract the pre-images for all actions and then group them by the action type. Next, we find the intersection of all pre-images for the same action to be the precondition. Similarly, for generating postconditions, we start by extracting the post-images for all actions and grouping them. The intersection of the post-images for each action is then calculated. This intersection is compared to the corresponding precondition, and any unchanged predicates are removed. The resulting set of predicates is returned as the postcondition. Collecting all actions with preconditions and postconditions, the planning domain is automatically written into a PDDL file, which is ready for task and motion planning.
Figure 4: Comparison of the success rates of basic planning tasks for increasing numbers of objects to baselines.
Figure 5: Comparison of the number of queries to motion planners during domain optimization to baselines.
@@ IV-D Domain Optimization by Generate-and-Test Search
The domain optimization process is based on the generate-and-test search method, a heuristic search technique that involves backtracking [Mahmood2013-ye]. Starting from $\mathbf{D_{\mathrm{top}}}$, the optimization process iterates through all possible domains until the domain successfully solves the validation problem set $\mathbf{Q}_{v}$ with the minimal number of predicates and actions. There are two assumptions made to perform the generate and test search:
@@
1:$\mathbf{D_{\mathrm{perturb}}}=\mathbf{D_{\mathrm{top}}}$triangleright Initialization.
2:$is\_solved=plan(\mathbf{D_{\mathrm{perturb}}},\mathbf{Q}_{v})$
3:$element\_list=\omega_{\mathrm{\mathbf{D}}}^{\mathrm{top}}$
4$\triangleright$ $top(\mathbf{\mathit{L}}$) returns the highest-scored element from $\mathbf{\mathit{L}}$.
5:while not $is\_solved$ do
@@
11:$\mathbf{D_{\mathrm{optm}}}=\mathbf{D_{\mathrm{expanded}}}=\mathbf{D_{\mathrm{perturb}}}$
@@
The optimization process aims to add or remove predicates (actions) from the initial guess until the optimal domain is found. The domain $\mathbf{D_{\mathrm{perturb}}}$ is initialized as $\mathbf{D_{\mathrm{top}}}$. There are two major stages in the optimization process: expansion and contraction.
@@ V Experiments
In this section, we introduce the baselines and experimental tasks. Through our experiments, we aim to answer the following questions: (Q1) How effective is planning using the inferred planning domain? (Q2) Can we generalize to unseen and more complex tasks? (Q3) During the testing phase, how much does our method reduce computational costs in domain generation?
@@ V-A Planning Baselines
We compared the planning success rate of our method to behavior cloning baselines to highlight its improved efficiency and generalizability. All baselines are trained per task type. The graph attention network training is implemented in PyTorch [paszke2017automatic] and PyTorch Geometric [Fey2019-lv].
BC-logical: Inspired by prior work [Zhu2021-sx], this baseline uses logical scene graphs to train a GAT task planning policy. The scene graph formulation is the same as that of our estimator. Using the current and goal logical scene graph as input, BC-logical outputs the next logical action to be executed.
@@
BC-hybrid: Inspired by recent papers combining continuous and logical data for imitation learning in TAMP [Khodeir2023-so, Lin2022-tc], BC-hybrid deviates from BC-logical by using both continuous and logical states of objects in the node features.
@@ V-B Domain Optimization Baselines
We also compare our domain optimization method to a few baselines to evaluate how much our method reduces computational cost.
@@
Blind hill-climbing: Inspired by the recent work in automatic planning domain generation [Kumar2023-nk, Silver2023-mi], this search method starts with an empty initial planning domain and searches through all possible planning domains via hill-climbing. This method searches for the planning domain with the highest planning success rate while maintaining the minimal planning domain size.
@@ V-C Experimental Tasks
Experiments are developed on the PyBullet simulation [benelot2018]. The task planner is based on the PDDLStream library [Garrett2021-ka], and the motion planner is based on the pybullet-planning library [Garrett2015-wc]. We designed nine basic tasks and three composed tasks, including classic cube manipulation, everyday task scenarios, and puzzle solving, to evaluate our method.
Basic tasks: In stacking, the robot needs to stack cubes into a tower in a specific order. In the unstacking task, the objective is to disassemble the tower of cubes. In sorting, the robot needs to cluster cubes into two groups at different locations. In washing, the robot needs to wash food ingredients in the kitchen sink. In grilling, the robot needs to grill food on the stove. The cooking task involves a combination of washing and grilling food. The table cleaning task indicates collecting objects from the table and storing them in a bin. The painting task involves drawing random figures on an object or a piece of paper. The last basic task is to solve tower of Hanoi.
@@ V-D Training and Testing Dataset
The training dataset includes 30 demonstrations per basic task (no composed task), each involving a solvable TAMP problem containing two to four objects. For baseline training, the solution trajectory for each TAMP problem is formulated as state-action pairs. The test dataset includes 870 different TAMP problems, encompassing both basic and composed tasks, with object counts in the planning environment ranging from two to nine. For each object count, 10 test problems are generated with randomized object positions, and each problem must be solved within 30 seconds.
@@ V-E Real Robot Experiments
To demonstrate the effectiveness of our system with real robots, we implemented two real-robot experiments: one involving the Tower of Hanoi and the other focusing on stacking multiple cube towers. We adopt NVIDIA’s FoundationPose [Wen_2024_CVPR] to capture visual demonstrations, which provides the 6D pose trajectories of all objects in the video. These poses are then consumed downstream to compute predicates concerning the objects in the scene. The real-robot validation problems are set up manually, with photos of the initial and goal configurations processed in the same way. The inferred planning domain is combined with the ROS Moveit motion planner and deployed on a Franka Research 3 arm.
@@
VI Results and Discussion
To address (Q1), we evaluate our method on various tasks and compare it to a few baselines. As illustrated in Figure 4, our method consistently outperforms the baselines in planning success rate across all basic tasks. While both our method and the baselines perform well on simpler tasks (painting, table cleaning, and sorting), the baselines degrade significantly on more complex tasks with larger decision spaces. To address (Q2), we evaluate our methods with planning problems that involve an increased number of objects, as well as unseen composed tasks. Figure 4 shows that our method maintains success rates above 90% even with up to nine objects, while the behavior cloning baseline method experiences a sharp decline as the number of objects increases. These results indicate that our method has a significant advantage in generalizing to a more complex experiment setup. Furthermore, as shown in Table I, our method demonstrates excellent planning performance on generalizing to unseen composed tasks. With just one human demonstration per task, our method successfully infers the planning domain, enabling efficient planning even when the number of objects increases and their positions are shuffled. Since we only have one trajectory per composed task, we did not compare our method with the baselines. Training a behavior cloning policy with just one trajectory would be trivial. To answer (Q3), we compare the computational cost of domain optimization to the baselines. The computational cost is assessed based on the number of queries made to the motion planner, which is the most time-intensive component of the optimization process. This metric is chosen because it provides a consistent comparison independent of the computer’s performance. As shown in Figure 5, our method drastically lowers the computational cost in the optimization process. The reduction in computational cost is more apparent for basic task types and attenuates when dealing with composed task types because the most relevant domain prediction becomes less accurate. Figure 6 shows our method’s testing planning success rate of the domain inferred with different sizes of the validation problem set $\mathbf{Q}_{v}$ via Algorithm 1. The results indicate that our method requires only a small $\mathbf{Q}_{v}$. For many tasks, the correct planning domain can be inferred with just one or two validation problems. Even in more challenging scenarios, five validation problems are sufficient. The real robot experiment, presented in Figure 7, demonstrates the effectiveness of our method on real robots. Remarkably, we only need to demonstrate once for the real robot to learn and solve similar TAMP problems.
Figure 7: Real robot experiment. Left: Towers of Hanoi. Right: Stack multiple towers

### v3

@@ IV-C Precondition and Postcondition Generation
The precondition is determined by identifying the intersections of the pre-images of the same action across multiple instances. We first extract the pre-images for all actions and then group them by the action type. Next, we find the intersection of all pre-images for the same action to be the precondition. Similarly, for generating postconditions, we start by extracting the post-images for all actions and grouping them. The intersection of the post-images for each action is then calculated. This intersection is compared to the corresponding precondition, and any unchanged predicates are removed. The resulting set of predicates is returned as the postcondition.
Figure 4: Success rates of basic planning tasks with increasing objects, compared to baselines (one demonstration per task).
Figure 5: Comparison of motion planner queries for domain optimization, ensuring all methods find the optimal domain.
@@ IV-D Domain Optimization by Generate-and-Test Search
The domain optimization process is based on the generate-and-test search method, a heuristic search technique that involves backtracking [21]. Starting from $\mathbf{D_{\mathrm{top}}}$, the optimization process iterates through all possible domains until the domain successfully solves the validation problem set $\mathbf{Q}_{v}$ with the minimal number of predicates and actions. There are two assumptions made for the search:
@@
[…]
5:while not is_solvedis\_solved do
@@
[…]
@@
The optimization process aims to add or remove predicates (actions) from the initial guess until the optimal domain is found. The domain $\mathbf{D_{\mathrm{perturb}}}$ is initialized as $\mathbf{D_{\mathrm{top}}}$. There are two major stages: expansion and contraction.
@@ V Experiments
In this section, we introduce the baselines and experimental tasks. We aim to answer the following questions: (Q1) How effective is planning using the inferred planning domain? (Q2) Can we generalize to more complex and unseen tasks? (Q3) During testing, how much does our method reduce computational costs of domain generation?
@@ V-A Planning Baselines
We compared the planning success rate of our method to behaviour cloning baselines to highlight its improved efficiency and generalizability. All baselines are trained per task type. The graph attention network training is implemented in PyTorch [27] and PyTorch Geometric [6].
BC-logical: Inspired by prior work [37], this baseline uses logical scene graphs to train a GAT task planning policy. The scene graph formulation is the same as that of our estimator. Using the current and goal logical scene graph as input, BC-logical outputs the next logical action to be executed.
@@
BC-hybrid: Inspired by recent papers combining continuous and logical data for imitation learning in TAMP [14, 18], BC-hybrid deviates from BC-logical by using both continuous and logical states of objects in the node features.
@@ V-B Domain Optimization Baselines
We compare our domain optimization method to several baselines to assess the potential reduction in computational cost for easier tasks, as we learns planning domains for simple tasks that can be used in complex ones.
@@
Blind hill-climbing: Inspired by the recent work in automatic planning domain generation [16, 31], this search method starts with an empty initial planning domain and searches through all possible planning domains via hill-climbing. This method searches for the planning domain with the highest planning success rate while maintaining the minimal planning domain size.
@@ V-C Experimental Tasks
Experiments are developed on the PyBullet simulation [5]. The task planner is based on the PDDLStream library [8], and the motion planner is based on the pybullet-planning library [7]. We designed nine basic tasks and three composed tasks to evaluate generalizability to unseen environments, where the tasks can be described by the same action and predicate set, with varying object counts and positions.
Basic tasks: In stacking, the robot needs to stack cubes into a tower in a specific order. In the unstacking task, the objective is to disassemble the tower of cubes. In sorting, the robot needs to cluster cubes into two groups at different locations. In washing, the robot needs to wash food ingredients in the kitchen sink. In grilling, the robot needs to grill food on the stove. The cooking task involves a combination of washing and grilling food. The table cleaning task indicates collecting objects from the table and storing them in a bin. The painting task involves drawing random figures on a piece of paper. The last basic task is to solve Towers of Hanoi.
@@ V-D Training and Testing Dataset
The training dataset includes 30 demonstrations per basic task, each involving a solvable TAMP problem containing two to four objects. The solution trajectory and the human demonstration are provided together as state-action pairs for baseline training. The test dataset includes 870 different TAMP problems, encompassing basic and composed tasks, with object counts in the planning environment ranging from two to nine. All planning problems are generated by uniformly sampling random object poses across the entire workspace (table). For each object count, 10 test problems are generated with randomized object positions, and each problem must be solved within 30 seconds.
@@ V-E Real Robot Experiments
To demonstrate the effectiveness of our system with real robots, we implemented two real-robot experiments: one involving the Tower of Hanoi and the other focusing on stacking multiple cube towers. We adopt NVIDIA’s FoundationPose [33] to capture visual demonstrations, which provide the 6D pose trajectories of all objects. These poses are then consumed downstream to compute predicates concerning the objects in the scene. The real-robot validation problems are set up manually, with photos of the initial and goal configurations processed in the same way. The inferred planning domain is combined with the ROS Moveit motion planner and deployed on a Franka Research 3 robot.
Figure 7: Real robot experiment. Left: Towers of Hanoi. Right: Stack towers
VI Results and Discussion
To address (Q1), we evaluate our method on various tasks and compare it to a few baselines. As illustrated in Figure 4, our method consistently outperforms the baselines in planning success rate across all basic tasks. While both our method and the baselines perform well on simpler tasks (painting, table cleaning, and sorting), the baselines degrade significantly on more complex tasks with larger decision spaces. Our baseline evaluation aligns with prior studies, reinforcing that behaviour cloning requires more comprehensive training data for generalization [16, 31].
To address (Q2), we evaluate our methods with planning problems that involve an increased number of objects, as well as unseen composed tasks. Figure 4 shows that our method maintains success rates above 90% even with up to nine objects, while the behaviour cloning baseline method experiences a sharp decline as the number of objects increases. These results indicate that our method has a significant advantage in generalizing to a more complex experiment setup. Furthermore, as shown in Table I, our method demonstrates excellent planning performance on generalizing to unseen composed tasks. With just one human demonstration per task, our method successfully infers the planning domain, enabling efficient planning even when the number of objects increases and their positions are shuffled. Since we only have one trajectory per composed task, we did not compare our method with the baselines. Training a behaviour cloning policy with just one trajectory would be trivial.
To answer (Q3), we compare the computational cost of domain optimization to the baselines. The computational cost is assessed based on the number of queries made to the motion planner, which is the most time-intensive component of the optimization process. This metric is chosen because it provides a consistent comparison independent of the computer’s performance. As shown in Figure 5, our method drastically lowers the computational cost in the optimization process. The reduction in computational cost is more apparent for basic task types and attenuates when dealing with composed task types because the most relevant domain prediction becomes less accurate. Figure 6 shows our method’s testing planning success rate of the domain inferred with different sizes of the validation problem set $\mathbf{Q}_{v}$ via Algorithm 1. The results indicate that our method requires only a small $\mathbf{Q}_{v}$. For many tasks, the correct planning domain can be inferred with just one or two validation problems. Even in more challenging scenarios, five validation problems are sufficient.
@@ VI Results and Discussion
The real robot experiment, presented in Figure 7, demonstrates the effectiveness of our method on a real-world system. Remarkably, we only need to demonstrate once for the real robot to learn and solve similar TAMP problems.